OpenStudy (wade123):

Find the derivative of f(x) = -9/x at x = 6

OpenStudy (wade123):

is a 1/4 correct?

OpenStudy (wade123):

i will give a medal

OpenStudy (anonymous):

the function can be written as -9 x^(-1) f'(x)= 9 x^(-2) so you get 9/6^2=1/2^2=1/4 yes

OpenStudy (wade123):

Find the derivative of f(x) = 6/x at x = -2. can you pleasee help me with this one? @sweetsunray

OpenStudy (wade123):

is it 3/2?